### v1

# Pushing Trade-Off Boundaries: Compact yet Effective Remote Sensing Change Detection

Abstract
Remote sensing change detection is essential for monitoring urban expansion, disaster assessment, and resource management, offering timely, accurate, and large-scale insights into dynamic landscape transformations. While deep learning has revolutionized change detection, the increasing complexity and computational demands of modern models have not necessarily translated into significant accuracy gains. Instead of following this trend, this study explores a more efficient approach, focusing on lightweight models that maintain high accuracy while minimizing resource consumption, which is an essential requirement for on-satellite processing. To this end, we propose FlickCD, which means quick flick then get great results, pushing the boundaries of the performance-resource trade-off. FlickCD introduces an Enhanced Difference Module (EDM) to amplify critical feature differences between temporal phases while suppressing irrelevant variations such as lighting and weather changes, thereby reducing computational costs in the subsequent change decoder. Additionally, the FlickCD decoder incorporates Local-Global Fusion Blocks, leveraging Shifted Window Self-Attention (SWSA) and Enhanced Global Self-Attention (EGSA) to efficiently capture semantic information at multiple scales, preserving both coarse- and fine-grained changes. Extensive experiments on four benchmark datasets demonstrate that FlickCD reduces computational and storage overheads by more than an order of magnitude while achieving state-of-the-art (SOTA) performance or incurring only a minor (<1% F1) accuracy trade-off. The implementation code is publicly available at https://github.com/xulsh8/FlickCD.
Keywords Change Detection, Remote Sensing, Lightweight Deep Learning
1 Introduction
With the rapid advancement of remote sensing technologies, it has become possible to acquire an increasing number of high-resolution remote sensing images. These abundant and high-quality data sources have greatly enhanced the development of various remote sensing applications, among which change detection (CD) has particularly benefited as a key task. The target of CD is to identify the changes in the same region by analyzing multi-temporal remote sensing images acquired at different times. This technique can be applied to various fields, such as disaster damage assessment [1, 2], urban planning [3, 4], agricultural management [5] and environment monitoring [6]. The CD techniques makes it possible to gain a more comprehensive understanding of the regional differences, enabling the prediction of future changes and the formulation of corresponding measures.
With the growing success of deep learning in various computer vision tasks, it has also emerged as a powerful tool for CD. However, deep learning-based CD methods face two major challenges. The first challenge is how to better balance model performance and resource requirements. To achieve better feature extraction capabilities, change detection models often increase their depth or width a lot. For example, RCTNet [7] increases depth by stacking convolutional layers to fuse multi-scale feature maps, while ChangeMamba [8] expands the receptive field using state-space models to capture global semantic information. However, these operations require a substantial number of parameters and extensive computational resources, making it challenging to deploy them on resource-constrained devices, such as drones and satellites, which are commonly used for acquiring remote sensing images. Currently, there are also researches on lightweight change detection models [9, 10]. Nevertheless, achieving lightweight design often results in performance degradation and leads to a reduced ability to extract difference information.
The second challenge is how to ignore irrelevant differences while focusing on capturing significant changes. Irrelevant differences can arise due to various factors, including lighting variations, atmospheric cloud conditions, seasonal changes and so on. To extract the critical changed information between two temporal images, various change detection models employ different approaches. Most models obtain difference information by computing the Manhattan distance between corresponding pixels in two temporal images [11, 9, 7]. Some models merge the two temporal feature maps along the channel dimension and use convolution operations to extract difference information [12, 10], while others perform layer-wise or multi-level feature fusion to achieve this [8, 13]. However, each of these methods has its own limitations: some of them fail to filter out irrelevant changes, making it difficult for the model to distinguish between insignificant differences and crucial changes. Others require substantial computational resources and a large number of parameters to obtain difference information, thereby increasing the overall resource demand of the model.
Figure 1: Comparison of parameters and relative performance between FlickCD(Ours) and others. The relative performance is calculated by normalizing the F1 scores across the four datasets, with the lowest value set to 0 and the highest value set to 1.
To address the aforementioned challenges, we propose FlickCD, a lightweight model that achieves an excellent balance between resource efficiency and performance. As shown in Figure 1, we select several lightweight models and state-of-the-art models for comparison. It can be observed that while maintaining a lightweight design, FlickCD achieves an average performance close to that of state-of-the-art models.
FlickCD adopts an encoder-decoder architecture, with RepViT [14] chosen as its encoder. This encoder has the advantages of low latency and a small number of parameters. Unlike most change detection models, we utilize only the feature maps extracted from the first three stages of the encoder, as we observed that incorporating higher-level features does not lead to significant performance improvement, while substantially increasing the number of parameters and computational cost. After extracting feature maps from two temporal images using a shared-weight encoder, we employ the Enhanced Difference Module(EDM) to capture the difference information between them. The lightweight EDM can enhance the difference features while suppressing the irrelevant noise, thereby helping the decoder more accurately extract the changed features of interest. Then, we design the decoder using a bottom-up, layer-by-layer fusion approach. Before fusion, we incorporate a Local-Global Fusion Block(LGFB) to further integrate features. LGFB primarily utilizes attention mechanisms and consist of two components: Sliding-Window Self-attention(SWSA) and Effective Global Self-attention(EGSA). SWSA applies windows to confine attention to local areas, enabling the model to focus more on the relationship between nearby pixels and improving the accuracy of local object recognition. Subsequently, EGSA is employed to consider the relationship between pixels across the global scope, which is particularly crucial when the changed regions consist of a lot of similar small areas. After multiscale extraction and fusion, the process yields a high-quality binary change map as the final result.
Through extensive experiment, we validate the effectiveness of the modules in FlickCD and demonstrate that FlickCD achieves a better trade-off and higher efficiency compared to other models. Our contributions can be summarized as follows:
  • We develop the lightweight change detection model FlickCD and design an efficient lightweight decoder capable of extracting both local and global semantic information effectively.
  • We propose the Enhanced Difference Module (EDM), which filters out irrelevant change noise while preserving critical difference information, thereby helping the model accurately capture change information.
  • We conducted experiments to demonstrate the efficiency of FlickCD among current change detection models. It achieves low parameter and computation costs while maintaining high performance, even achieves state-of-the-art on one dataset.
2 Related Work
2.1 Deep Learning-based Change Detection
With the remarkable success and high performance of deep learning techniques, the change detection (CD) community has increasingly adopted them to address the CD task. One of the earliest CNN-based approaches is the FC series [15], which introduced two fundamental strategies for handling dual-temporal inputs: (1) early fusion, where the two input images are first concatenated and then fed into the network; and (2) Siamese architectures, which process the two images with shared-weight encoders followed by a difference computation. These two strategies have laid the foundation for most subsequent CD models. After that, the emergence of Transformer architectures, which leverage attention mechanisms, has brought new possibilities to CD. For instance, BIT [11] adopts a CNN backbone for encoding and decoding, while employing a Transformer to capture differences between the input images. ChangeFormer [12] further advances this direction by directly applying a shared-weight Transformer to handle the CD task. In addition to the conventional global self-attention, various attention mechanisms tailored for CD have also been explored. These include spatial and channel attention [16, 17, 18], as well as cross-attention mechanism [19], which enhance the effectiveness and robustness of CD models.
2.2 Lightweight Strategies in Deep Learning
Striking a perfect balance between model efficiency and performance is always a critical research direction. To enable the development of deep learning models on resource-constrained devices, a variety of lightweight strategies have been proposed across different model architectures. In convolutional neural networks(CNNs)-based strategies, depthwise separable convolution [20] extracts features by decoupling the spatial and channel dimensions. Group convolution [21] divides channels into groups and performing convolution operations within each group independently. More recently, FasterNet [22] introduce partial convolution, which compute only a subset of channels in the feature maps. In Transformer-based strategies, they are primarily implemented in two ways. First, limit the attention computation scope to local windows or sparse regions [14, 23, 24]. Although this strategy may result in information loss, allocating limited parameters to more likely relevant tokens can lead to minimal performance degradation while reducing redundant information. Second, utilize weight-sharing mapping to obtain the attention matrix [23, 25]. Applying the same linear mapping to different regions with correlations can reduce the number of parameters and map them to the same high-dimensional space. As for the CD tasks, there has also been research on lightweight models. USSFC-Net [9] uses dilated convolutions cyclically as the basic module, and construct the Encoder and Decoder by stacking this module. TinyCD [10] employs the three-stage EfficientNet [26] as its backbone and utilizes group convolution to compute difference information. Although these networks have a low number of parameters, they suffer from high latency and lower performance.
Figure 2: (a) Overall architecture of the FlickCD model. (b) Structural illustration of the Enhanced Difference Module (EDM) and Fusion module.
3 Method
3.1 Overview
The overview of the model structure is shown in the left side of Figure 2. At first, input the bitemporal images, $T_{1}$ and $T_{2}$, both with dimensions $W\times H\times 3$, into the weight-sharing encoder. We selected the lightest version of RepViT [14] with pre-trained weights as the encoder. Then, we only use the feature maps from the first three stages of the encoder. On one hand, our experiments demonstrate that the feature map from the fourth stage does not improve the model’s performance. On the other hand, removing the fourth stage can reduce the number of parameters and computational complexity. This process can be expressed as:
$f_{1\_1},f_{1\_2},f_{1\_3}=Encoder(T_{1}),$
$f_{2\_1},f_{2\_2},f_{2\_3}=Encoder(T_{2}).$
The encoder extracts features hierarchically, where downsampling and channel doubling are applied at each stage. For the feature maps $f_{1\_j}$ and $f_{2\_j},j=(1,2,3)$, the spatial resolution at each stage is $\frac{W}{2^{j+1}}\times\frac{H}{2^{j+1}}$. Then, we use the Enhanced Difference Module(EDM) to extract the difference feature between $f_{1}$ and $f_{2}$ at each stage. Additionally, to facilitate the multi-stage feature fusion in the decoder, the channel dimension of the feature maps from all three stages are unified. We formulate this process as:
$D_{i}=EDM_{i}(f_{1\_i},f_{2\_i}),\{i=1,2,3\}.$
After that, we propose a bottom-up lightweight decoder, which mainly consists of the Local-Global Fusion Block(LGFB) and additional Effective Global Self-attention(EGSA). Through layer-by-layer fusion and upsampling, it generates the final binary result map.
3.2 Enhanced Difference Module
How to extract the difference information between two temporal images effectively is one of the main challenges of the CD task. A common approach for extracting the difference is to directly subtract the two feature maps and take the absolute value, which corresponds to the Manhattan distance. However, since this approach computes pixel-wise differences with equal weighting across all pixels, trivial differences caused by irrelevant factors are preserved alongside meaningful change information and passed to the decoder. This makes it difficult for the decoder to accurately identify critical differences, leading to a decrease in accuracy. Moreover, additional computational resources and parameters may be required to effectively distinguish between relevant and irrelevant information. To address this issue, we propose the Enhanced Difference Module(EDM), which dynamically assigns weights to the pixel-wise differences. This enables the model to filter out irrelevant difference and enhance the identification of critical change information.
The detailed architecture of the EDM is illustrated on the right side of Fig 2. For a given stage $EDM_{i}$, the inputs are the corresponding temporal feature maps $f_{1\_i}$ and $f_{2\_i}$. Since the magnitude of pixel-wise differences should take into account local contextual information, we first apply a depthwise separable convolution to enhance local feature aggregation. This is followed by a Squeeze-and-Excitation block and a $1\times 1$ convolution layer, which are used to perform dynamic channel-wise weighting and adjust the number of output channels accordingly.
After the above operations, we obtain the transformed feature maps $f^{\prime}_{1\_i}$ and $f^{\prime}_{2\_i}$, which are then fed into the Get Difference Attention module to generate the difference feature map $D_{i}$. This module is inspired by the core concept of attention mechanisms: measuring the similarity between vectors via dot products to assign corresponding token weights. Specifically, we first apply identical linear projection to $f^{\prime}_{1\_i}$ and $f^{\prime}_{2\_i}$, mapping them into a shared high-dimensional space and obtain the query Q and the key K. Next, we compute the dot product between corresponding vectors in Q and K to derive the initial attention mask M. Unlike standard attention mechanisms, our goal is to emphasize the features that exhibit greater differences. To this end, we invert the attention mask M, such that vector pairs with larger dot products (i.e., similar directions) receive lower weights, while those with smaller dot products (i.e., more distinct features) are assigned higher weights.This inversion encourages the model to focus more on significantly different features. Finally, we apply a Sigmoid function to normalize the inverted scores, resulting in the final attention mask $M^{\prime}$. To mitigate the risk of vanishing gradients caused by excessively large dot product values, a scaling factor $\sqrt{d_{k}}$(where $d_{k}$ represents the vector dimension) is applied during the dot product computation. This process can be expressed as:
$\textbf{Q}=Conv_{1\times 1}(f^{\prime}_{1\_i}),$
$\textbf{K}=Conv_{1\times 1}(f^{\prime}_{2\_i}),$
$\textbf{M}\supset m_{ij}=\frac{q_{ij}\cdot k_{ij}}{\sqrt{d_{k}}},\{i=1,2,\dots,W;j=1,2,\dots,H\},$
$^{\prime}=Sigmoid(-\textbf{M}).$ M ) .
Finally, we compute teh value matrix V by calculating the Manhattan distance between the transformed feature maps $f^{\prime}_{1\_i}$ and $f^{\prime}_{2\_i}$. The final output $D_{i}$ is then obtained by Hadamard product of the attention mask $\textbf{M}^{\prime}$ and value matrix V.
$\textbf{V}=Conv^{\prime}_{1\times 1}(|f^{\prime}_{1\_i}-f^{\prime}_{2\_i}|),$
$D_{i}=\textbf{M}^{\prime}\otimes\textbf{V}.$
Figure 3: (a)Structural diagram of the Sliding Window Self-Attention (SWSA) module. (b)Detailed architecture of the Local-Global Fusion Block (LGFB). (c)Structural diagram of the Efficient Global Self-Attention (EGSA) module.
3.3 Local-Global Fusion Block
After obtaining the difference map $D_{i}$ between the two temporal feature maps, it is essential to further capture the relational information among the changed regions. To this end, we extract relationship features at both local and global scales, motivated by the following observations:
  • On the local scale, when a specific region is surronded by other changed areas, it is highly likely that this region has also undergone change. Modeling such local relationships helps mitigate the presence of hols or fragmentation within detected changed regions.
  • On the global scale, some images contain multiple instances of similar changes, for example, repeated structural changes in buildings. Capturing global contextual relationships enables the model to identify and enhance the consistency among these semantically similar change patterns, thereby improving the extraction accuracy for such repeated changes across the image.
To achieve the aforementioned objectives, we propose the Local-Global Fusion Blick(LGFB). This module consists of two key components: Sliding-Window Self-Attention(SWSA) and Efficient Global Self-Attention(EGSA). The former is designed to capture local relational information among pixels within spatially adjacent regions, while the latter focuses on modeling long-range dependencies and similarities among changed regions across the entire image. The detailed architecture of LGFB is illustrated in Figure 3. Structurally, LGFB is a Transformer-like module. However, conventional self-attention mechanisms typically incur substantial computational and parametric overhead. To address this, we adopt a more efficient design, which is described in detail below.
Token Mixer & Channel Mixer. Inspired by the RepViT [14], we decompose the attention computation into two stages: a token mixer and a channel mixer, where the former is responsible for aggregating spatial information and the latter focuses on inter-channel interaction. Specifically, in both the SWSA and EGSA modules, we adopt depthwise separable convolutions(DWConv) for linear projections. While this projection lacks explicit interaction across channels, our subsequent experiments demonstrate that this design substantially reduces the number of parameters without compromising model performance. The resulting projected matrices are then fed into the attention mechanism, enabling spatial information exchange with minimal parameter and computational overhead—serving as the token mixer. This process can be formally expressed as:
$Q=DWConv_{1}(I),$
$K=DWConv_{2}(I),$
$V=DWConv_{3}(I),$
$O=Sigmoid(\frac{QK^{T}}{\sqrt{d}})V+I.$
Here, $I$ represents the input to the SWSA and EGSA modules. Subsequently, a multilayer perceptron (MLP) composed of two $1\times 1$ convolutional layers is employed to perform information exchange along the channel dimension, serving as the channel mixer. This process can be formally expressed as:
$O^{\prime}=MLP(BN(O))+O.$
Sliding Window Self-Attention.To enable the model to focus on local pixel-level relationships, we constrain the attention computation within a sliding window. A stride parameter $s$ is introduced to control how the window moves across the entire feature map, producing the final attention output. Both the window size $w$ and the stride $s$ are treated as hyperparameters. When the window size is equal to the stride ($w=s$), the feature map is partitioned into non-overlapping patches, and attention is computed independently within each patch. In contrast, when the window size is larger than the stride ($w>s$), only a portion of the computed attention within each window is retained as a patch, resulting in overlapping receptive fields where the patch size is $s\times s$. In scenarios where change regions exhibit complex shapes, using a larger window with overlapping receptive fields enhances spatial continuity across neighboring regions. Conversely, for regions with low shape complexity, such as regular building changes, setting the window size equal to the stride is sufficient to capture local patterns. It avoids unnecessary contextual redundancy while reducing additional computational overhead.
Efficient Global Self-Attention.To capture global relationships among change regions and further mitigate the spatial discontinuity introduced by patch-wise processing in SWSA, we introduce Efficient Global Self-Attention (EGSA) following SWSA. In EGSA, we first apply linear projections to the input $D_{L}$ to obtain the query $Q_{G}$, key $K_{G}$, and value $V_{G}$ matrices. The $Q_{G}$ matrix retains the same spatial resolution as the input to preserve detailed spatial information, while the $K_{G}$ and $V_{G}$ matrices are downsampled to a lower resolution. This downsampling serves to compress global information, effectively reducing both the parameter count and computational cost. Specifically, the $K_{G}$ and $V_{G}$ matrices are obtained via convolution operations, where both the kernel size and stride are set to match the patch size used in SWSA. This design ensures that each token in $K_{G}$ and $V_{G}$ corresponds to a patch from the input $D_{L}$. We then compute the attention weights between each pixel in $Q_{G}$ and all patch-level tokens in $K_{G}$, enabling the model to aggregate long-range contextual information. The final output feature map is generated by applying the computed attention weights to the $V_{G}$ matrix.
Hierarchical Feature Integration.The lightweight decoder takes as input the multi-scale difference features $D_{i}(i=1,2,3)$ generated from three hierarchical stages. Among them, higher-level difference maps exhibit lower spatial resolution but encode richer semantic information about change regions, while lower-level difference maps preserve finer spatial details such as object boundaries but tend to be less reliable in distinguishing actual changes. To leverage the complementary advantages of different levels, we adopt a progressive fusion strategy. As illustrated on the left side of Figure 2, each higher-level difference map is processed by the LGFB, and then fused with the corresponding lower-level difference map via element-wise addition. A depthwise separable convolution is then applied to refine the fused features, achieving a local adjustment. However, this local fusion alone is not sufficient to capture long-range dependencies across spatially distant regions. Therefore, we further incorporate EGSA after the local fusion to perform global refinement and enhance the fusion quality.
Figure 4: Changed region characteristics in four datasets. For each dataset, samples are divided by the number of change regions ($\leq$4 on the left, $>$4 on the right). The x-axis denotes shape complexity (perimeter/area); the y-axis shows area ratio for the left plots and region similarity (variance of shape complexity) for the right plots.
4 Experiment and Discussion
4.1 Datasets and Experiment Settings
Dataset Preparation and Analysis. To validate the effectiveness of FlickCD, we conduct experiments on four widely-used change detection datasets: SYSU [27], WHU [28], LEVIR+ [29], and CDD [30]. To better interpret the experimental results, we perform a statistical analysis on the characteristics of change regions within each dataset, as shown in Figure 4. We categorize the samples in each dataset into two groups based on the number of change regions: few-region samples and many-region samples, using a threshold of 4 change regions. For few-region samples, we analyze the area proportion of change regions and their shape complexity. For many-region samples, we focus on the inter-region dissimilarity and shape complexity. The analysis reveals that:
  • SYSU primarily consists of samples with a small number of change regions, where most changes occupy a large area and exhibit relatively high shape complexity.
  • CDD predominantly contains samples with many change regions, characterized by high variability in inter-region similarity and complex region shapes.
  • WHU and LEVIR+ focus mainly on building-related changes, leading to a large number of multi-region samples. In these datasets, the change regions tend to exhibit high inter-region similarity and low shape complexity, reflecting the relatively regular structure of architectural changes.
Experiment Settings. To accommodate the varying characteristics of change regions across datasets, we configure the hyperparameters of the Sliding-Window Self-Attention (SWSA) module, namely the window size $w$ and stride $s$ individually for each dataset. Given that higher-level feature maps have lower spatial resolution, we assign smaller window sizes at deeper layers accordingly.For the SYSU and CDD datasets, the change regions tend to exhibit higher shape complexity, requiring a larger receptive field to improve the accuracy of feature extraction. Accordingly, we set the window sizes to 8,8,16 with corresponding strides of 4,4,8 for these datasets. In contrast, the change regions in WHU and LEVIR+ are generally more regular and structurally simple. For these datasets, we adopt an equal window size and stride configuration to reduce computational overhead without sacrificing performance. Specifically, we set the parameters to 4,4,8 for WHU and 4,8,8 for LEVIR+. For model training, the batch size is set to 32. The learning rate is set to 0.0005 for all datasets except WHU, which uses a lower learning rate of 0.0002. We train the model for 100 epochs on the SYSU and WHU datasets, 200 epochs on LEVIR+, and 250 epochs on CDD.
Table 1: Comparative Analysis of model complexity and performance on SYSU and WHU datasets. Bold indicates the best result, and underline denotes the second-best. All values are presented as percentages(%).
Methods Param. FLOPs(G) SYSU WHU
Rec. Pre. OA F1 IoU Rec. Pre. OA F1 IoU
FC-EF[15] 1.35M 6.23 77.92 78.94 89.89 78.43 64.51 73.10 87.89 99.10 79.81 66.41
FC-Siam-Diff[15] 1.35M 8.51 79.05 80.03 90.41 79.54 66.03 85.06 90.02 99.41 87.47 77.73
FC-Siam-Conc[15] 1.55M 9.72 79.25 79.48 90.28 79.36 65.79 86.75 90.82 99.46 88.74 79.76
TinyCD[10] 285.12K 2.81 80.08 80.76 90.80 80.42 67.25 87.30 91.93 99.50 89.55 81.09
USSFC-Net[9] 1.52M 6.32 79.23 78.18 89.89 78.70 64.88 90.15 92.75 99.59 91.43 84.21
ChangeFormer[12] 41.03M 276.81 75.61 82.61 90.49 78.95 65.23 90.44 94.09 99.62 92.23 85.58
RCTNet[7] 13.94M 23.99 78.64 84.79 91.64 81.60 68.92 91.58 95.09 99.68 93.30 87.44
ChangeMamba[8] 53.99M 61.69 78.61 88.27 92.49 83.16 71.17 93.58 95.57 99.74 94.56 89.69
BiFA[31] 9.87M 105.75 80.43 87.52 92.68 83.83 72.16 91.37 95.50 99.68 93.39 87.60
FlickCD(Ours) 1.89M 4.16 80.28 88.03 92.77 83.97 72.38 93.18 96.51 99.75 94.81 90.14
Table 2: The model complexity and performance comparison of different models on CDD and LEVIR+ datasets.
Method Param. FLOPs(G) CDD LEVIR+
Rec. Pre. OA F1 IoU Rec. Pre. OA F1 IoU
FC-EF[15] 1.35M 6.23 91.46 93.94 98.13 92.68 86.37 74.42 74.04 97.89 74.23 59.02
FC-Siam-Diff[15] 1.35M 8.51 94.44 95.91 98.76 95.17 90.79 80.05 80.28 98.39 80.17 66.90
FC-Siam-Conc[15] 1.55M 9.72 93.04 95.20 98.49 94.11 88.87 80.92 81.93 98.50 81.42 68.67
TinyCD[10] 285.12K 2.81 93.40 94.73 98.48 94.06 88.79 82.23 82.86 98.58 82.54 70.28
USSFC-Net[9] 1.52M 6.32 96.57 93.99 98.87 95.26 90.95 77.19 80.46 98.31 78.79 65.00
ChangeFormer[12] 41.03M 276.81 94.51 94.82 98.74 94.67 89.87 81.45 84.14 98.61 82.77 70.61
RCTNet[7] 13.94M 23.99 96.99 97.80 99.33 97.39 94.92 84.58 85.09 98.77 84.83 73.66
ChangeMamba[8] 53.99M 61.69 98.21 97.66 99.51 97.93 95.95 86.99 87.70 98.97 87.34 77.53
BiFA[31] 9.87M 105.75 95.63 95.64 98.97 95.63 91.63 84.44 81.93 98.61 83.16 71.18
FlickCD(Ours) 1.89M 4.16 97.09 98.18 99.39 97.63 95.37 84.59 88.09 98.91 86.30 75.90
Figure 5: Visualization results on the WHU and LEVIR+ dataset, where white represents true positive, black represents true negative, green represents false positive, and red represents false negative.
4.2 Overall Comparison
Comparison Baselines. To demonstrate the trade-off advantage of FlickCD in change detection tasks, we compare it against both lightweight and high-performance CD models. The lightweight models include the FC series [15] (FC-EF, FC-Siam-Diff, FC-Siam-Conc), TinyCD [10], and USSFC-Net [9], all of which are based on convolutional neural networks (CNNs). The high-performance models include RCTNet [7](CNN-based), ChangeFormer [12] and BiFA [31](Transformer-based), and ChangeMamba [8](based on the recent Mamba architecture). Through this comparison, we aim to highlight the efficiency advantage of FlickCD over high-performance models, and its performance advantage over lightweight models.
Quantitative Results. We primarily compare the F1 scores of different models across four datasets, as this metric provides a comprehensive evaluation of model performance. As shown in Table 1 and 2, it is evident that models with fewer parameters than FlickCD perform significantly worse across all datasets. Compared with the models that have more parameters, FlickCD achieves the highest performance in SYSU and WHU datasets, and maintains performance on par with the best-performing model ChangeMamba in CDD and LEVIR+ datasets. Specifically, FlickCD achieves an F1-score of 83.97% on the SYSU dataset and 94.81% on the WHU dataset, surpassing the second-best model by 0.14% and 0.25%, respectively. Although FlickCD does not outperform ChangeMamba in terms of F1 score on the CDD and LEVIR+ datasets, it maintains a comparable performance level and ranks second among the selected models.
Qualitative Results. To gain a more intuitive and detailed understanding of model performance in CD tasks, we select samples from the SYSU, WHU, and LEVIR+ datasets and compare the output of five resource-intensive models with FlickCD. As shown in Figures 5 and 6, FlickCD demonstrates superior recognition performance on challenging change regions. Specifically, for Figure 6 and sample (3) in Figures 5, some color variations caused by illumination differences are actually irrelevant changes. However, the comparison models mistakenly classify them as significant change areas, while FlickCD correctly identifies and filters out these irrelevant differences. Furthermore, although FlickCD achieves a slightly lower F1-score than ChangeMamba on the LEVIR+ dataset, it shows better extraction performance in cases involving multiple similar change regions, such as sample (4) in Figure 5. This visual comparison highlights FlickCD’s stronger ability to distinguish irrelevant differences and more accurate boundary extraction for change regions.
Model Efficiency Analysis To compare the computational efficiency and resource requirements of different models, we report the number of parameters, FLOPs, throughput, and inference time for each model. Specifically, the inference time refers to the time required to process a single sample, while the throughput is calculated by dividing the number of samples in one epoch by the total runtime of that epoch. The corresponding results are presented in Table 1 and 3. The FC series, TinyCD, and USSFC-Net significantly reduce the parameter count at the cost of considerable performance degradation. However, FC models and USSFC-Net still incur relatively high computational costs. Among the high-performance models, although they achieve competitive results across various datasets, their resource demands are substantial. For example, ChangeMamba, which demonstrates a good balance of accuracy across datasets, requires 53.99M parameters and 31.3 ms inference time per sample. In contrast, FlickCD achieves comparable performance with only 1.89M parameters,approximately 3.5% of ChangeMamba, and a lower inference time of 9.2 ms, which is less than one-third of that of ChangeMamba. In addition, FlickCD achieves a throughput that is 2 to 3 times higher than that of the high-performance models. These results clearly demonstrate the efficiency of FlickCD in terms of parameter count, computation cost, and latency, while maintaining strong performance.
Table 3: Comparison of Throughput and Inference Time across different models . Higher Throughput and Lower Inference Time Indicate Better Performance.
Methods Throughput(samples/s) Inference Time (ms)
FC-EF[15] 402.82 1.5
FC-Siam-Diff[15] 323.10 2.1
FC-Siam-Conc[15] 326.00 2.1
TinyCD[10] 166.81 5.0
USSFC-Net[9] 68.66 13.4
ChangeFormer[12] 40.42 11.4
RCTNet[7] 56.03 16.7
ChangeMamba[8] 30.82 31.3
BiFA[31] 32.32 19.8
FlickCD (Ours) 95.08 9.2
4.3 Ablation Study
Effectiveness of EDM To validate the effectiveness of the Enhanced Difference Module(EDM) in FlickCD, we first conducted a visual ablation study, with the results shown in Figure 6. On the SYSU dataset, FlickCD exhibits a stronger ability to distinguish between irrelevant and meaningful changes. To confirm that this capability stems from the EDM, we remove EDM and retrained the model, using a direct Manhattan distance to compute the difference features between the two temporal inputs. As illustrated in the results, the model fails to distinguish irrelevant changes from significant ones without EDM, similar to other baseline methods. In addition to the visual comparison, we also performed quantitative ablation experiments across all datasets, as shown in Table 4. By integrating the EDM into the base architecture, we observe consistent improvements in F1-scores across all datasets, with the most significant gain of 0.65% on SYSU. These findings from both visual and quantitative perspectives demonstrate the effectiveness of the EDM module in enhancing the model’s discriminative capability.
Figure 6: Visualization results on the SYSU dataset, showcasing the outputs of other models, FlickCD with EDM, and FlickCD without EDM. The comparison highlights FlickCD’s superior difference discrimination capability and validates the effectiveness of the EDM.
Table 4: Ablation Study of Key Modules. All values are presented as percentages(%).
Module F1
EDM SWSA EGSA SYSU WHU CDD LEVIR+
80.41 93.34 96.60 85.91
81.06 93.98 96.77 86.19
82.62 93.73 97.18 85.63
82.66 93.89 97.33 85.94
83.97 94.81 97.63 86.30
Effectiveness of SWSA and EGSA. To further verify the effectiveness of Sliding-Window Self-Attention(SWSA) and Efficient Global Self-Attention(EGSA) in analyzing difference information, we conducted ablation experiments involving these two modules on all datasets. The results are summarized in Table 4, revealing two distinct trends:
  • For the SYSU and CDD datasets, introducing either module individually leads to performance improvements, and combining both results in further gains.
  • In contrast, on the WHU and LEVIR+ datasets, using only one of the modules causes a performance drop, while combining both modules yields a notable performance increase.
These findings, when correlated with the dataset characteristics illustrated in Figure 4, provide further insights. Both WHU and LEVIR+ datasets contain a large number of multi-instance change regions, with high similarity among regions and low shape complexity. In such cases, omitting one attention scale (local or global) leads to a greater loss than the gain achieved by using only one. This suggests that both local and global contextual analyses are essential for effectively capturing meaningful change information in datasets characterized by numerous, similar, and relatively simple change regions.
Table 5: Comparison of the parameter reduction and performance impact of two lightweight strategies, w/o dpConv indicates that depthwise separable convolutions are not used in SWSA and EGSA. Four stages refers to using outputs from all four stages of the backbone.
Strategy Param. F1
SYSU WHU CDD LEVIR+
Base 1.89M 83.97 94.81 97.63 86.30
w/o dpConv 3.60M 83.51 94.05 97.54 85.68
Four stages 3.19M 82.67 94.05 96.99 85.99
Lightweight Strategy. In FlickCD, we investigated two lightweight design strategies that potentially introduce some information loss but significantly reduce the model size:
  • Employing depthwise separable convolutions for linear projections in both EGSA and SWSA.
  • Using feature maps from only the first three encoder stages for change detection.
To assess the potential trade-offs introduced by these strategies, we conducted ablation experiments, with the results summarized in Table 5. The findings indicate that both strategies reduce the model’s parameter count by nearly half, yet no performance degradation was observed across the four datasets. This suggests that the information loss induced by these strategies is non-critical. Specifically, while depthwise separable convolutions discard inter-channel interactions, this limitation is compensated by the subsequent MLP module, which facilitates cross-channel communication. Furthermore, the third encoder stage already provides sufficient semantic abstraction for change region extraction, and incorporating the fourth stage does not further improve performance.
5 Conclusion
We propose a lightweight model, FlickCD, which maintains high performance while minimizing resource requirements. The model is primarily composed of convolutional operations and attention mechanisms. Specifically, we design the Enhanced Difference Module (EDM) to enhance bitemporal image difference information while reducing irrelevant noise. Additionally, we introduce two lightweight feature extraction modules, SWSA and EGSA, which effectively extract the local and global features while maintaining efficiency. To evaluate the effectiveness of FlickCD, we conducted extensive experiments on four datasets, analyzing its resource consumption and inference speed, along with ablation studies. The results consistently demonstrate our model’s low computational cost and high performance. This research provides new insights and methodologies for deploying change detection models on mobile devices in the future.

### v2

@@
 FlickCD introduces an Enhanced Difference Module (EDM) to amplify critical feature differences between temporal phases while suppressing irrelevant variations such as lighting and weather changes, thereby reducing computational costs in the subsequent change decoder.
-Additionally, the FlickCD decoder incorporates Local-Global Fusion Blocks, leveraging Shifted Window Self-Attention (SWSA) and Enhanced Global Self-Attention (EGSA) to efficiently capture semantic information at multiple scales, preserving both coarse- and fine-grained changes.
+Additionally, the FlickCD decoder incorporates Local-Global Fusion Blocks, leveraging Shifted Window Self-Attention (SWSA) and Efficient Global Self-Attention (EGSA) to effectively capture semantic information at multiple scales, preserving both coarse- and fine-grained changes.
 Extensive experiments on four benchmark datasets demonstrate that FlickCD reduces computational and storage overheads by more than an order of magnitude while achieving state-of-the-art (SOTA) performance or incurring only a minor (<1% F1) accuracy trade-off.
 The implementation code is publicly available at https://github.com/xulsh8/FlickCD.
-Keywords Change Detection, Remote Sensing, Lightweight Deep Learning
+Keywords Remote Sensing; Change Detection; Lightweight Deep Learning
 1 Introduction
@@
 This technique can be applied to various fields, such as disaster damage assessment [1, 2], urban planning [3, 4], agricultural management [5] and environment monitoring [6].
-The CD techniques makes it possible to gain a more comprehensive understanding of the regional differences, enabling the prediction of future changes and the formulation of corresponding measures.
+The CD techniques make it possible to gain a more comprehensive understanding of the regional differences, enabling the prediction of future changes and the formulation of corresponding measures.
 With the growing success of deep learning in various computer vision tasks, it has also emerged as a powerful tool for CD.
@@
 However, these operations require a substantial number of parameters and extensive computational resources, making it challenging to deploy them on resource-constrained devices, such as drones and satellites, which are commonly used for acquiring remote sensing images.
-Currently, there are also researches on lightweight change detection models [9, 10].
-Nevertheless, achieving lightweight design often results in performance degradation and leads to a reduced ability to extract difference information.
+Currently, several studies have also been conducted on lightweight change detection models [9, 10].
+Nevertheless, achieving a lightweight design often results in performance degradation and leads to a reduced ability to extract difference information.
 The second challenge is how to ignore irrelevant differences while focusing on capturing significant changes.
-Irrelevant differences can arise due to various factors, including lighting variations, atmospheric cloud conditions, seasonal changes and so on.
+Irrelevant differences can arise due to various factors, including lighting variations, atmospheric conditions, and seasonal changes.
 To extract the critical changed information between two temporal images, various change detection models employ different approaches.
-Most models obtain difference information by computing the Manhattan distance between corresponding pixels in two temporal images [11, 9, 7].
+Most models obtain difference information by subtracting the feature maps and taking the absolute value.
+[11, 9, 7].
 Some models merge the two temporal feature maps along the channel dimension and use convolution operations to extract difference information [12, 10], while others perform layer-wise or multi-level feature fusion to achieve this [8, 13].
-However, each of these methods has its own limitations: some of them fail to filter out irrelevant changes, making it difficult for the model to distinguish between insignificant differences and crucial changes.
+However, each of these methods has its own limitations: some of them fail to filter out irrelevant changes, making it difficult for the model to distinguish between insignificant differences and critical changes.
 Others require substantial computational resources and a large number of parameters to obtain difference information, thereby increasing the overall resource demand of the model.
@@
 To address the aforementioned challenges, we propose FlickCD, a lightweight model that achieves an excellent balance between resource efficiency and performance.
-As shown in Figure 1, we select several lightweight models and state-of-the-art models for comparison.
+As shown in Fig.
+1, we select several lightweight models and state-of-the-art models for comparison.
 It can be observed that while maintaining a lightweight design, FlickCD achieves an average performance close to that of state-of-the-art models.
@@
 Before fusion, we incorporate a Local-Global Fusion Block (LGFB) to further integrate features.
-LGFB primarily utilizes attention mechanisms and consist of two components: Sliding-Window Self-attention(SWSA) and Effective Global Self-attention(EGSA).
-SWSA applies windows to confine attention to local areas, enabling the model to focus more on the relationship between nearby pixels and improving the accuracy of local object recognition.
-Subsequently, EGSA is employed to consider the relationship between pixels across the global scope, which is particularly crucial when the changed regions consist of a lot of similar small areas.
+LGFB primarily utilizes attention mechanisms and consists of two components: Sliding-Window Self-Attention (SWSA) and Efficient Global Self-Attention (EGSA).
+SWSA uses window-based attention to constrain focus to local regions, which helps the model better capture spatial relationships between nearby pixels and improves the accuracy of local object recognition.
+Subsequently, EGSA is employed to consider the relationship between pixels across the global scope, which is particularly critical when the changed regions consist of a lot of similar small areas.
 After multiscale extraction and fusion, the process yields a high-quality binary change map as the final result.
-Through extensive experiment, we validate the effectiveness of the modules in FlickCD and demonstrate that FlickCD achieves a better trade-off and higher efficiency compared to other models.
+Through extensive experiments, we validate the effectiveness of the modules in FlickCD and demonstrate that FlickCD achieves a better trade-off and higher efficiency compared to other models.
 Our contributions can be summarized as follows:
@@
   • We propose the Enhanced Difference Module (EDM), which filters out irrelevant change noise while preserving critical difference information, thereby helping the model accurately capture change information.
-  • We conducted experiments to demonstrate the efficiency of FlickCD among current change detection models.
-It achieves low parameter and computation costs while maintaining high performance, even achieves state-of-the-art on one dataset.
+  • We conduct experiments to validate the efficiency of FlickCD among current CD models.
+It achieves low parameter count and computation cost while maintaining high performance, and even achieves state-of-the-art on some datasets.
 2 Related Work
@@
 These two strategies have laid the foundation for most subsequent CD models.
-After that, the emergence of Transformer architectures, which leverage attention mechanisms, has brought new possibilities to CD.
+After that, the emergence of Transformer architectures, which leverage the attention mechanisms, has brought new possibilities to CD.
 For instance, BIT [11] adopts a CNN backbone for encoding and decoding, while employing a Transformer to capture differences between the input images.
@@
 In addition to the conventional global self-attention, various attention mechanisms tailored for CD have also been explored.
-These include spatial and channel attention [16, 17, 18], as well as cross-attention mechanism [19], which enhance the effectiveness and robustness of CD models.
+These include spatial and channel attention [16, 17, 18], as well as the cross-attention mechanism [19], which enhance the effectiveness and robustness of CD models.
 2.2 Lightweight Strategies in Deep Learning
@@
 To enable the development of deep learning models on resource-constrained devices, a variety of lightweight strategies have been proposed across different model architectures.
-In convolutional neural networks(CNNs)-based strategies, depthwise separable convolution [20] extracts features by decoupling the spatial and channel dimensions.
-Group convolution [21] divides channels into groups and performing convolution operations within each group independently.
-More recently, FasterNet [22] introduce partial convolution, which compute only a subset of channels in the feature maps.
+In CNN-based strategies, Depth-wise Separable Convolution [20] extracts features by decoupling the spatial and channel dimensions.
+Group Convolution [21] divides channels into groups and performs convolution operations within each group independently.
+More recently, FasterNet [22] introduces Partial Convolution, which computes only a subset of channels in the feature maps.
 In Transformer-based strategies, they are primarily implemented in two ways.
@@
 Second, utilize weight-sharing mapping to obtain the attention matrix [23, 25].
-Applying the same linear mapping to different regions with correlations can reduce the number of parameters and map them to the same high-dimensional space.
-As for the CD tasks, there has also been research on lightweight models.
-USSFC-Net [9] uses dilated convolutions cyclically as the basic module, and construct the Encoder and Decoder by stacking this module.
+Applying the same linear mapping to correlated regions reduces parameter count and projects them into the same high-dimensional space.
+There has also been research on lightweight models for CD tasks.
+USSFC-Net [9] uses dilated convolutions cyclically as the basic module, and constructs the Encoder and Decoder by stacking this module.
 TinyCD [10] employs the three-stage EfficientNet [26] as its backbone and utilizes group convolution to compute difference information.
@@
 3.1 Overview
-The overview of the model structure is shown in the left side of Figure 2.
-At first, input the bitemporal images, $T_{1}$ and $T_{2}$, both with dimensions $W\times H\times 3$, into the weight-sharing encoder.
+The overview of the model structure is shown in the Fig.
+2(a).
+At first, input the bi-temporal images, $\mathbf{T}_{1}$ and $\mathbf{T}_{2}$, both with dimensions $W\times H\times 3$, into the weight-sharing encoder.
 We selected the lightest version of RepViT [14] with pre-trained weights as the encoder.
@@
 This process can be expressed as:
-$f_{1\_1},f_{1\_2},f_{1\_3}=Encoder(T_{1}),$
-$f_{2\_1},f_{2\_2},f_{2\_3}=Encoder(T_{2}).$
+$f_{1\_1},f_{1\_2},f_{1\_3}=Encoder(\mathbf{T}_{1}),$
+$f_{2\_1},f_{2\_2},f_{2\_3}=Encoder(\mathbf{T}_{2}).$
 The encoder extracts features hierarchically, where downsampling and channel doubling are applied at each stage.
-For the feature maps $f_{1\_j}$ and $f_{2\_j},j=(1,2,3)$, the spatial resolution at each stage is $\frac{W}{2^{j+1}}\times\frac{H}{2^{j+1}}$.
+For the feature maps $f_{1\_i}$ and $f_{2\_i},i\in\{1,2,3\}$, the spatial resolution at each stage is $\frac{W}{2^{i+1}}\times\frac{H}{2^{i+1}}$.
 Then, we use the Enhanced Difference Module (EDM) to extract the difference feature between $f_{1}$ and $f_{2}$ at each stage.
@@
 We formulate this process as:
-$D_{i}=EDM_{i}(f_{1\_i},f_{2\_i}),\{i=1,2,3\}.$
-After that, we propose a bottom-up lightweight decoder, which mainly consists of the Local-Global Fusion Block(LGFB) and additional Effective Global Self-attention(EGSA).
+$D_{i}=EDM_{i}(f_{1\_i},f_{2\_i}),i\in\{1,2,3\}.$
+After that, we propose a bottom-up lightweight decoder, which mainly consists of the Local-Global Fusion Block (LGFB) and additional Efficient Global Self-attention (EGSA).
 Through layer-by-layer fusion and upsampling, it generates the final binary result map.
@@
 How to extract the difference information between two temporal images effectively is one of the main challenges of the CD task.
-A common approach for extracting the difference is to directly subtract the two feature maps and take the absolute value, which corresponds to the Manhattan distance.
-However, since this approach computes pixel-wise differences with equal weighting across all pixels, trivial differences caused by irrelevant factors are preserved alongside meaningful change information and passed to the decoder.
+A common approach for extracting the difference is to directly subtract the two feature maps and take the absolute value.
+However, since this approach computes pixel-wise differences with equal weights across all pixels, differences caused by irrelevant factors are preserved alongside meaningful change information and passed to the decoder.
 This makes it difficult for the decoder to accurately identify critical differences, leading to a decrease in accuracy.
-Moreover, additional computational resources and parameters may be required to effectively distinguish between relevant and irrelevant information.
 To address this issue, we propose the Enhanced Difference Module (EDM), which dynamically assigns weights to the pixel-wise differences.
-This enables the model to filter out irrelevant difference and enhance the identification of critical change information.
-The detailed architecture of the EDM is illustrated on the right side of Fig 2.
-For a given stage $EDM_{i}$, the inputs are the corresponding temporal feature maps $f_{1\_i}$ and $f_{2\_i}$.
-Since the magnitude of pixel-wise differences should take into account local contextual information, we first apply a depthwise separable convolution to enhance local feature aggregation.
+This enables the model to filter out irrelevant differences and enhance the identification of critical change information.
+The detailed architecture of the EDM is illustrated in Fig.
+2(b).
+For a given stage $EDM_{i}$, the inputs are the corresponding feature maps $f_{1\_i}$ and $f_{2\_i}$.
+Since the magnitude of pixel-wise differences should take local contextual information into account, we first apply a depthwise separable convolution to enhance local feature aggregation.
 This is followed by a Squeeze-and-Excitation block and a $1\times 1$ convolution layer, which are used to perform dynamic channel-wise weighting and adjust the number of output channels accordingly.
-After the above operations, we obtain the transformed feature maps $f^{\prime}_{1\_i}$ and $f^{\prime}_{2\_i}$, which are then fed into the Get Difference Attention module to generate the difference feature map $D_{i}$.
+After the above operations, we obtain the feature maps $f^{\prime}_{1\_i}$ and $f^{\prime}_{2\_i}$, which are then fed into the Difference-Aware Attention module to generate the difference feature map $D_{i}$.
 This module is inspired by the core concept of attention mechanisms: measuring the similarity between vectors via dot products to assign corresponding token weights.
@@
 Unlike standard attention mechanisms, our goal is to emphasize the features that exhibit greater differences.
-To this end, we invert the attention mask M, such that vector pairs with larger dot products (i.e., similar directions) receive lower weights, while those with smaller dot products (i.e., more distinct features) are assigned higher weights.This inversion encourages the model to focus more on significantly different features.
-Finally, we apply a Sigmoid function to normalize the inverted scores, resulting in the final attention mask $M^{\prime}$.
+To this end, we invert the attention mask M, such that vector pairs with larger dot products (i.e., similar directions) receive lower weights, while those with smaller dot products (i.e., more distinct features) are assigned higher weights.
+This inversion encourages the model to focus more on significantly different features.
+Finally, we apply a Sigmoid function to normalize the inverted scores, resulting in the final attention mask $\textbf{M}^{\prime}$.
 To mitigate the risk of vanishing gradients caused by excessively large dot product values, a scaling factor $\sqrt{d_{k}}$ (where $d_{k}$ represents the vector dimension) is applied during the dot product computation.
 This process can be expressed as:
-$\textbf{Q}=Conv_{1\times 1}(f^{\prime}_{1\_i}),$
-$\textbf{K}=Conv_{1\times 1}(f^{\prime}_{2\_i}),$
-$\textbf{M}\supset m_{ij}=\frac{q_{ij}\cdot k_{ij}}{\sqrt{d_{k}}},\{i=1,2,\dots,W;j=1,2,\dots,H\},$
-$^{\prime}=Sigmoid(-\textbf{M}).$ M ) .
-Finally, we compute teh value matrix V by calculating the Manhattan distance between the transformed feature maps $f^{\prime}_{1\_i}$ and $f^{\prime}_{2\_i}$.
-The final output $D_{i}$ is then obtained by Hadamard product of the attention mask $\textbf{M}^{\prime}$ and value matrix V.
+$\textbf{Q}=Conv_{1\times 1}(f^{\prime}_{1\_i}),\textbf{K}=Conv_{1\times 1}(f^{\prime}_{2\_i}),$
+$\textbf{M}\supset m_{ij}=\frac{q_{ij}\cdot k_{ij}}{\sqrt{d_{k}}},i\in\{1,2,\dots,W\};j\in\{1,2,\dots,H\},$
+$\textbf{M}^{\prime}=Sigmoid(-\textbf{M}).$
+Finally, we compute the value matrix V by taking the element-wise absolute difference between the feature maps $f^{\prime}_{1\_i}$ and $f^{\prime}_{2\_i}$.
+The final output $D_{i}$ is then obtained by the Hadamard product of the attention mask $\textbf{M}^{\prime}$ and value matrix V.
 $\textbf{V}=Conv^{\prime}_{1\times 1}(|f^{\prime}_{1\_i}-f^{\prime}_{2\_i}|),$
-$D_{i}=\textbf{M}^{\prime}\otimes\textbf{V}.$
+$D_{i}=\textbf{M}^{\prime}\odot\textbf{V}.$
 Figure 3: (a)Structural diagram of the Sliding Window Self-Attention (SWSA) module.
@@
 3.3 Local-Global Fusion Block
-After obtaining the difference map $D_{i}$ between the two temporal feature maps, it is essential to further capture the relational information among the changed regions.
+After obtaining the difference map $D_{i}$ between the two feature maps, it is essential to further capture the relational information among the changed regions.
 To this end, we extract relationship features at both local and global scales, motivated by the following observations:
-  • On the local scale, when a specific region is surronded by other changed areas, it is highly likely that this region has also undergone change.
-Modeling such local relationships helps mitigate the presence of hols or fragmentation within detected changed regions.
-  • On the global scale, some images contain multiple instances of similar changes, for example, repeated structural changes in buildings.
-Capturing global contextual relationships enables the model to identify and enhance the consistency among these semantically similar change patterns, thereby improving the extraction accuracy for such repeated changes across the image.
-To achieve the aforementioned objectives, we propose the Local-Global Fusion Blick(LGFB).
-This module consists of two key components: Sliding-Window Self-Attention(SWSA) and Efficient Global Self-Attention(EGSA).
-The former is designed to capture local relational information among pixels within spatially adjacent regions, while the latter focuses on modeling long-range dependencies and similarities among changed regions across the entire image.
-The detailed architecture of LGFB is illustrated in Figure 3.
+  • On the local scale, when a specific region is surrounded by other changed areas, it is highly likely that this region has also undergone change.
+Modeling such local relationships helps mitigate the presence of holes or fragmentation within detected changed regions.
+  • On the global scale, by leveraging global contextual information, the model can enhance overall feature representations, thereby increasing the discriminability between changed and unchanged regions and facilitating more accurate delineation of change areas.
+To achieve the aforementioned objectives, we propose the Local-Global Fusion Block (LGFB).
+This module consists of two key components: Sliding Window Self-Attention (SWSA) and Efficient Global Self-Attention (EGSA).
+The former is designed to capture local relational information among pixels within spatially adjacent regions, while the latter focuses on modeling long-range dependencies and enhancing feature representation.
+The detailed architecture of LGFB is illustrated in Fig.
+3.
 Structurally, LGFB is a Transformer-like module.
-However, conventional self-attention mechanisms typically incur substantial computational and parametric overhead.
+However, traditional self-attention mechanisms are often computationally and parameter-intensive.
 To address this, we adopt a more efficient design, which is described in detail below.
@@
 Inspired by the RepViT [14], we decompose the attention computation into two stages: a token mixer and a channel mixer, where the former is responsible for aggregating spatial information and the latter focuses on inter-channel interaction.
-Specifically, in both the SWSA and EGSA modules, we adopt depthwise separable convolutions(DWConv) for linear projections.
-While this projection lacks explicit interaction across channels, our subsequent experiments demonstrate that this design substantially reduces the number of parameters without compromising model performance.
-The resulting projected matrices are then fed into the attention mechanism, enabling spatial information exchange with minimal parameter and computational overhead—serving as the token mixer.
+Specifically, in both the SWSA and EGSA modules, we adopt Depth-wise Separable Convolution (DWConv) for linear projections.
+The projected matrices are then fed into the attention mechanism to efficiently exchange spatial information, acting as a lightweight token mixer.
 This process can be formally expressed as:
-$Q=DWConv_{1}(I),$
-$K=DWConv_{2}(I),$
-$V=DWConv_{3}(I),$
+$Q,K,V=DWConv_{1}(I),DWConv_{2}(I),DWConv_{3}(I),$
 $O=Sigmoid(\frac{QK^{T}}{\sqrt{d}})V+I.$
 Here, $I$ represents the input to the SWSA and EGSA modules.
-Subsequently, a multilayer perceptron (MLP) composed of two $1\times 1$ convolutional layers is employed to perform information exchange along the channel dimension, serving as the channel mixer.
+Subsequently, a multi-layer perceptron (MLP) composed of two $1\times 1$ convolutional layers is employed to perform information exchange along the channel dimension, serving as the channel mixer.
 This process can be formally expressed as:
@@
 In contrast, when the window size is larger than the stride ($w>s$), only a portion of the computed attention within each window is retained as a patch, resulting in overlapping receptive fields where the patch size is $s\times s$.
-In scenarios where change regions exhibit complex shapes, using a larger window with overlapping receptive fields enhances spatial continuity across neighboring regions.
+In scenarios where changed regions exhibit complex shapes, using a larger window with overlapping receptive fields enhances spatial continuity across neighboring regions.
 Conversely, for regions with low shape complexity, such as regular building changes, setting the window size equal to the stride is sufficient to capture local patterns.
 It avoids unnecessary contextual redundancy while reducing additional computational overhead.
-Efficient Global Self-Attention.To capture global relationships among change regions and further mitigate the spatial discontinuity introduced by patch-wise processing in SWSA, we introduce Efficient Global Self-Attention (EGSA) following SWSA.
+Efficient Global Self-Attention.To capture global relationships among changed regions and further mitigate the spatial discontinuity introduced by patch-wise processing in SWSA, we introduce Efficient Global Self-Attention (EGSA) following SWSA.
 In EGSA, we first apply linear projections to the input $D_{L}$ to obtain the query $Q_{G}$, key $K_{G}$, and value $V_{G}$ matrices.
 The $Q_{G}$ matrix retains the same spatial resolution as the input to preserve detailed spatial information, while the $K_{G}$ and $V_{G}$ matrices are downsampled to a lower resolution.
-This downsampling serves to compress global information, effectively reducing both the parameter count and computational cost.
+This downsampling serves to compress global information, effectively reducing computational cost.
 Specifically, the $K_{G}$ and $V_{G}$ matrices are obtained via convolution operations, where both the kernel size and stride are set to match the patch size used in SWSA.
@@
 The final output feature map is generated by applying the computed attention weights to the $V_{G}$ matrix.
-Hierarchical Feature Integration.The lightweight decoder takes as input the multi-scale difference features $D_{i}(i=1,2,3)$ generated from three hierarchical stages.
-Among them, higher-level difference maps exhibit lower spatial resolution but encode richer semantic information about change regions, while lower-level difference maps preserve finer spatial details such as object boundaries but tend to be less reliable in distinguishing actual changes.
+Hierarchical Feature Integration.The lightweight decoder takes the multi-scale difference features $D_{i},i\in\{1,2,3\}$ generated from three hierarchical stages, as input.
+Among them, higher-level difference maps have lower spatial resolution but capture richer semantic information about changed regions, whereas lower-level maps retain finer spatial details, such as object boundaries, but are less reliable in distinguishing actual changes.
 To leverage the complementary advantages of different levels, we adopt a progressive fusion strategy.
-As illustrated on the left side of Figure 2, each higher-level difference map is processed by the LGFB, and then fused with the corresponding lower-level difference map via element-wise addition.
-A depthwise separable convolution is then applied to refine the fused features, achieving a local adjustment.
+As illustrated in Fig.
+2(a), each higher-level difference map is processed by the LGFB, and then fused with the corresponding lower-level difference map via element-wise addition.
+A Depth-wise Separable Convolution is then applied to refine the fused features, achieving a local adjustment.
 However, this local fusion alone is not sufficient to capture long-range dependencies across spatially distant regions.
@@
 Figure 4: Changed region characteristics in four datasets.
-For each dataset, samples are divided by the number of change regions ($\leq$4 on the left, $>$4 on the right).
+For each dataset, samples are divided by the number of changed regions ($\leq$4 on the left, >>4 on the right).
 The x-axis denotes shape complexity (perimeter/area); the y-axis shows area ratio for the left plots and region similarity (variance of shape complexity) for the right plots.
@@
 4.1 Datasets and Experiment Settings
-Dataset Preparation and Analysis.
 To validate the effectiveness of FlickCD, we conduct experiments on four widely-used change detection datasets: SYSU [27], WHU [28], LEVIR+ [29], and CDD [30].
-To better interpret the experimental results, we perform a statistical analysis on the characteristics of change regions within each dataset, as shown in Figure 4.
-We categorize the samples in each dataset into two groups based on the number of change regions: few-region samples and many-region samples, using a threshold of 4 change regions.
-For few-region samples, we analyze the area proportion of change regions and their shape complexity.
+To better interpret the experimental results, we perform a statistical analysis on the characteristics of changed regions within each dataset, as shown in Fig.
+4.
+We categorize the samples in each dataset into two groups based on the number of changed regions: few-region samples and many-region samples, using a threshold of 4 changed regions.
+For few-region samples, we analyze the area proportion of changed regions and their shape complexity.
 For many-region samples, we focus on the inter-region dissimilarity and shape complexity.
 The analysis reveals that:
-  • SYSU primarily consists of samples with a small number of change regions, where most changes occupy a large area and exhibit relatively high shape complexity.
-  • CDD predominantly contains samples with many change regions, characterized by high variability in inter-region similarity and complex region shapes.
+  • SYSU primarily consists of samples with a small number of changed regions, where most changes occupy a large area and exhibit relatively high shape complexity.
+  • CDD mainly consists of samples with numerous changed regions, which show large variation in similarity between regions and exhibit complex shapes.
   • WHU and LEVIR+ focus mainly on building-related changes, leading to a large number of multi-region samples.
-In these datasets, the change regions tend to exhibit high inter-region similarity and low shape complexity, reflecting the relatively regular structure of architectural changes.
+In these datasets, the changed regions tend to exhibit high inter-region similarity and low shape complexity, reflecting the relatively regular structure of architectural changes.
 Experiment Settings.
-To accommodate the varying characteristics of change regions across datasets, we configure the hyperparameters of the Sliding-Window Self-Attention (SWSA) module, namely the window size $w$ and stride $s$ individually for each dataset.
-Given that higher-level feature maps have lower spatial resolution, we assign smaller window sizes at deeper layers accordingly.For the SYSU and CDD datasets, the change regions tend to exhibit higher shape complexity, requiring a larger receptive field to improve the accuracy of feature extraction.
+To accommodate the varying characteristics of changed regions across datasets, we configure the hyperparameters of the Sliding-Window Self-Attention (SWSA) module, namely the window size $w$ and stride $s$ individually for each dataset.
+Given that higher-level feature maps have lower spatial resolution, we assign smaller window sizes at deeper layers accordingly.For the SYSU and CDD datasets, the changed regions tend to exhibit higher shape complexity, requiring a larger receptive field to improve the accuracy of feature extraction.
 Accordingly, we set the window sizes to 8,8,16 with corresponding strides of 4,4,8 for these datasets.
-In contrast, the change regions in WHU and LEVIR+ are generally more regular and structurally simple.
-For these datasets, we adopt an equal window size and stride configuration to reduce computational overhead without sacrificing performance.
+In contrast, the changed regions in WHU and LEVIR+ are generally more regular and structurally simple.
+For these datasets, we adopt smaller window sizes and strides configuration to reduce computation.
 Specifically, we set the parameters to 4,4,8 for WHU and 4,8,8 for LEVIR+.
@@
 FlickCD(Ours) 1.89M 4.16 97.09 98.18 99.39 97.63 95.37 84.59 88.09 98.91 86.30 75.90
-Figure 5: Visualization results on the WHU and LEVIR+ dataset, where white represents true positive, black represents true negative, green represents false positive, and red represents false negative.
 4.2 Overall Comparison
-Comparison Baselines.
-To demonstrate the trade-off advantage of FlickCD in change detection tasks, we compare it against both lightweight and high-performance CD models.
-The lightweight models include the FC series [15] (FC-EF, FC-Siam-Diff, FC-Siam-Conc), TinyCD [10], and USSFC-Net [9], all of which are based on convolutional neural networks (CNNs).
-The high-performance models include RCTNet [7](CNN-based), ChangeFormer [12] and BiFA [31](Transformer-based), and ChangeMamba [8](based on the recent Mamba architecture).
+To demonstrate the trade-off advantage of FlickCD in CD tasks, we compare it against both lightweight and high-performance CD models.
+The lightweight models include the FC series (FC-EF, FC-Siam-Diff, FC-Siam-Conc) [15], TinyCD [10], and USSFC-Net [9].
+The high-performance models include RCTNet [7], ChangeFormer [12], BiFA [31], and ChangeMamba [8].
 Through this comparison, we aim to highlight the efficiency advantage of FlickCD over high-performance models, and its performance advantage over lightweight models.
@@
 Compared with the models that have more parameters, FlickCD achieves the highest performance in SYSU and WHU datasets, and maintains performance on par with the best-performing model ChangeMamba in CDD and LEVIR+ datasets.
-Specifically, FlickCD achieves an F1-score of 83.97% on the SYSU dataset and 94.81% on the WHU dataset, surpassing the second-best model by 0.14% and 0.25%, respectively.
 Although FlickCD does not outperform ChangeMamba in terms of F1 score on the CDD and LEVIR+ datasets, it maintains a comparable performance level and ranks second among the selected models.
@@
 To gain a more intuitive and detailed understanding of model performance in CD tasks, we select samples from the SYSU, WHU, and LEVIR+ datasets and compare the output of five resource-intensive models with FlickCD.
-As shown in Figures 5 and 6, FlickCD demonstrates superior recognition performance on challenging change regions.
-Specifically, for Figure 6 and sample (3) in Figures 5, some color variations caused by illumination differences are actually irrelevant changes.
-However, the comparison models mistakenly classify them as significant change areas, while FlickCD correctly identifies and filters out these irrelevant differences.
-Furthermore, although FlickCD achieves a slightly lower F1-score than ChangeMamba on the LEVIR+ dataset, it shows better extraction performance in cases involving multiple similar change regions, such as sample (4) in Figure 5.
-This visual comparison highlights FlickCD’s stronger ability to distinguish irrelevant differences and more accurate boundary extraction for change regions.
+As shown in Fig.
+5 and 6, FlickCD demonstrates superior recognition performance on challenging changed regions.
+Specifically, for Fig.
+5(3) and Fig.
+6, some color variations caused by illumination differences are actually irrelevant changes.
+However, the comparison models mistakenly classify them as significant changes, while FlickCD identifies and filters out these irrelevant differences.
+Furthermore, although FlickCD achieves a slightly lower F1-score than ChangeMamba on the LEVIR+ dataset, it shows better extraction performance in cases involving multiple similar changed regions, such as Fig.
+5(4).
+This visual comparison highlights FlickCD’s stronger ability to distinguish irrelevant differences and more accurate extraction for changed regions.
+Figure 5: Visualization results on the WHU and LEVIR+ dataset, where white represents true positive, black represents true negative, green represents false positive, and red represents false negative.
 Model Efficiency Analysis To compare the computational efficiency and resource requirements of different models, we report the number of parameters, FLOPs, throughput, and inference time for each model.
@@
 The FC series, TinyCD, and USSFC-Net significantly reduce the parameter count at the cost of considerable performance degradation.
-However, FC models and USSFC-Net still incur relatively high computational costs.
+However, the FC series and USSFC-Net still have relatively high FLOPs values.
 Among the high-performance models, although they achieve competitive results across various datasets, their resource demands are substantial.
@@
 4.3 Ablation Study
-Effectiveness of EDM To validate the effectiveness of the Enhanced Difference Module(EDM) in FlickCD, we first conducted a visual ablation study, with the results shown in Figure 6.
-On the SYSU dataset, FlickCD exhibits a stronger ability to distinguish between irrelevant and meaningful changes.
-To confirm that this capability stems from the EDM, we remove EDM and retrained the model, using a direct Manhattan distance to compute the difference features between the two temporal inputs.
+Effectiveness of EDM To validate the effectiveness of the Enhanced Difference Module (EDM) in FlickCD, we first conducted a visual ablation study, with the results shown in Fig.
+6.
+On the SYSU dataset, FlickCD exhibits a stronger ability to distinguish between irrelevant and critical changes.
+To confirm that this capability stems from the EDM, we removed EDM and retrained the model, obtaining the difference features between the two inputs by subtracting their feature maps and taking the absolute value.
 As illustrated in the results, the model fails to distinguish irrelevant changes from significant ones without EDM, similar to other baseline methods.
@@
 By integrating the EDM into the base architecture, we observe consistent improvements in F1-scores across all datasets, with the most significant gain of 0.65% on SYSU.
-These findings from both visual and quantitative perspectives demonstrate the effectiveness of the EDM module in enhancing the model’s discriminative capability.
+These findings from both visual and quantitative perspectives demonstrate the effectiveness of the EDM in enhancing the model’s discriminative capability.
 Figure 6: Visualization results on the SYSU dataset, showcasing the outputs of other models, FlickCD with EDM, and FlickCD without EDM.
-The comparison highlights FlickCD’s superior difference discrimination capability and validates the effectiveness of the EDM.
 Table 4: Ablation Study of Key Modules. All values are presented as percentages(%).
@@
 Effectiveness of SWSA and EGSA.
-To further verify the effectiveness of Sliding-Window Self-Attention(SWSA) and Efficient Global Self-Attention(EGSA) in analyzing difference information, we conducted ablation experiments involving these two modules on all datasets.
+To verify the effectiveness of Sliding-Window Self-Attention (SWSA) and Efficient Global Self-Attention (EGSA) in analyzing difference information, we conducted ablation experiments involving these two modules on all datasets.
 The results are summarized in Table 4, revealing two distinct trends:
@@
   • In contrast, on the WHU and LEVIR+ datasets, using only one of the modules causes a performance drop, while combining both modules yields a notable performance increase.
-These findings, when correlated with the dataset characteristics illustrated in Figure 4, provide further insights.
-Both WHU and LEVIR+ datasets contain a large number of multi-instance change regions, with high similarity among regions and low shape complexity.
+These findings, when correlated with the dataset characteristics illustrated in Fig.
+4, provide further insights.
+Both WHU and LEVIR+ datasets contain a large number of multi-instance changed regions, with high similarity among regions and low shape complexity.
 In such cases, omitting one attention scale (local or global) leads to a greater loss than the gain achieved by using only one.
-This suggests that both local and global contextual analyses are essential for effectively capturing meaningful change information in datasets characterized by numerous, similar, and relatively simple change regions.
+This suggests that both local and global contextual analyses are essential for effectively capturing meaningful changed information in datasets characterized by numerous, similar, and relatively simple changed regions.
+Figure 7: Visualization of intermediate feature maps and attention heatmaps: (a) Difference map before EDM; (b) Attention heatmap of EDM; (c) Feature map after EDM; (d) Feature map after SWSA; (e) Feature map after EGSA; (f) Final result.
+Visualization of Intermediate Features.
+To further demonstrate the effectiveness of each module, we visualize several intermediate feature maps and attention heatmaps, as shown in Fig.
+7.
+The difference map before EDM processing fails to accurately detect the changed regions.
+In contrast, the attention heatmap generated by EDM reallocates the weights of different regions, resulting in a more precise segmentation of the changed areas (Fig.
+7(c)).
+Subsequently, SWSA refines the delineation of local changed regions, while EGSA further enhances the global difference information.
+Finally, the model outputs an accurate binary change detection map.
 Table 5: Comparison of the parameter reduction and performance impact of two lightweight strategies, w/o dpConv indicates that depthwise separable convolutions are not used in SWSA and EGSA.
@@
 Lightweight Strategy. In FlickCD, we investigated two lightweight design strategies that potentially introduce some information loss but significantly reduce the model size:
-  • Employing depthwise separable convolutions for linear projections in both EGSA and SWSA.
+  • Employing depth-wise separable convolutions for linear projections in both EGSA and SWSA.
   • Using feature maps from only the first three encoder stages for change detection.
@@
 This suggests that the information loss induced by these strategies is non-critical.
-Specifically, while depthwise separable convolutions discard inter-channel interactions, this limitation is compensated by the subsequent MLP module, which facilitates cross-channel communication.
-Furthermore, the third encoder stage already provides sufficient semantic abstraction for change region extraction, and incorporating the fourth stage does not further improve performance.
+Specifically, while depth-wise separable convolutions discard inter-channel interactions, this limitation is compensated by the subsequent MLP, which facilitates cross-channel communication.
+In addition, the third encoder stage already provides sufficient semantic information for changed region extraction, and incorporating the fourth stage does not further improve performance.
 5 Conclusion
 We propose a lightweight model, FlickCD, which maintains high performance while minimizing resource requirements.
-The model is primarily composed of convolutional operations and attention mechanisms.
-Specifically, we design the Enhanced Difference Module (EDM) to enhance bitemporal image difference information while reducing irrelevant noise.
+The model is primarily composed of convolution operations and attention mechanisms.
+Specifically, we design the Enhanced Difference Module (EDM) to enhance bi-temporal image difference information while reducing irrelevant noise.
 Additionally, we introduce two lightweight feature extraction modules, SWSA and EGSA, which effectively extract the local and global features while maintaining efficiency.
@@
 The results consistently demonstrate our model’s low computational cost and high performance.
-This research provides new insights and methodologies for deploying change detection models on mobile devices in the future.
+This research provides new insights and methodologies for deploying change detection models on resource-constrained devices in the future.